OpenStudy (anonymous):

Use the elimination method to find all solutions to the system of equations x^2-y^2=25 2X62-y^2=x+55 There are three solutions for each (x,y)

OpenStudy (anonymous):

add them together get \[3x^2=x+80\] solve \[3x^2-x-80=0\]